Broward Manufacturing recently reported the following information: Net income $515,000 ROA 12% Interest expense $195,700 Accounts payable and accruals $1,050,000 Broward's tax rate is 40%. Broward finances with only debt and common equity, so it has no preferred stock. 40% of its total invested capital is debt, while 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC). Round your answers to two decimal places. BEP % ROE % ROIC %
Answer #1

Net Income = $ 515000, Tax Rate = 40 %

Profit Before Tax (PBT) = Net Income / (1 - Tax Rate) = 515000 / (1-0.4) = $ 858333.33

Return on Asset (ROA) = (Net Income / Total Assets) x 100 = 12 %

(515000 / Total Assets) = 12 / 100

Total Assets = 515000 / 0.12 = $ 4291666.67

Any firm's asset needs to be financed by an equal corresponding amount of capital known as invested capital. Hence, in this case total firm asset = total invested capital.In this case, the firm's invested capital is entirely financed by common equity and debt. Therefore, 40 % of the firm's total assets are debt and the remaining is common equity.

Debt = 0.4 x 4291666.67 = $ 1716666.67

Common Equity = 0.6 x 4291666.67 = $ 2575000

Interest Expense = $ 195700

EBIT = PBT + Interest Expense = 858333.33 + 195700 = $ 1054033.33

BEP = EBIT / Total Assets = 1054033.33 / 4219666.67 ~ 0.2456 or 24.56 %

ROE = Net Income / Common Equity = 515000 / 2575000 ~ 0.2 or 20 %

ROCI = EBIT x (1-Tax Rate) / Total Invested Capital = 10504033.33 x (1-0.4) / 4291666.67 ~ 0.14736 or 14.74 %